FS Corp estimates that its demand function is as: Q = 620 – 12P + 20A + 10Y - 8P* where Q - quantity demanded per month, P - product’s price (in £), A- firm’s advertising expenditures (in £1000 per month), Y-per capita disposable income (in £1000), and P* - price of GT Corp.

a) During the next two years, per capita disposable income is expected to increase by £4,000 and GT is expected to increase its price by £6. Estimate the effect this will this have on FS’s sales volume, stating any necessary assumptions.

b) If FS wants to change its advertising by enough to offset the above effects, by how much must it do so

c) If FS’s current price is £50 and it spends £10,000 per month on promotion, while per capita income is £25,000 and GT’s price is £60, calculate the promotional elasticity of demand at the end of two years with the change in advertising.

d) What can be said about the relationship between the products of FS and GT


1
Expert's answer
2021-06-16T04:26:25-0400

a) If during the next two years, per capita disposable income is expected to increase by £4,000 and GT is expected to increase its price by £6, then FS’s sales volume will decrease by 4×10 - 8×6 = -8 units.

b) If FS wants to change its advertising by enough to offset the above effects, then it should increase it by 8/20 = 0.4 or £400.

c) If FS’s current price is £50 and it spends £10,000 per month on promotion, while per capita income is £25,000 and GT’s price is £60, then Q = 620 – 12×50 + 20×10 + 10×25 - 8×60 = -10.

The promotional elasticity of demand at the end of two years with the change in advertising is:

Ed = -12×50/(-10) = 60.

d) The products of FS and GT are complements.